OpenStudy (anonymous):

2. Solve using the substitution method. Show your work. If the system has no solution or an infinite number of solutions, state this. 8x – 5y = 3 7x + y = 51

OpenStudy (anas.p):

Lets take the second equation and make y the subject. \[y=51-7x\] Now place this value in the first equation. \[8x-5(51-7x)=3\]Now solve this equation. \[8x-255+35x=3\]\[43x=258\]\[x=258/43\]\[x=6\] Now place this equation in any of the above equations. Lets put it in the first equation.\[8(6)-5y=3\]\[48-5y=3\]\[48-3=5y\]\[45=5y\]\[45/5=y\]\[y=9\]